When rubidium ions are heated to a high temperature, two lines are observed in its line...

  1. When rubidium ions are heated to a high temperature, two lines are observed in its line spectrum at wavelengths (a) 7.67 x 10-7 m and (b) 3.99 x 10-7 m. What are the frequencies of the two lines? What color are the lines?
  2. Choose the answer that best completes the following: The existence of atomic line spectra indicates that. . .
    1. the energy of a photon depends on its frequency.
    2. all orbitals in s subshells must be spherical in shape.
    3. electrons in atoms orbit the nucleus in a circular path.
    4. electrons in atoms are restricted to certain energy levels.
    5. isotopes of an element contain different numbers of protons in the nucleus.
  3. How far from the nucleus in angstroms (1 angstrom = 1x10-10 m) is the electron in a hydrogen atom if it has an energy of -7.67x10-19 J?
  4. State the Heisenberg uncertainty principle. Briefly describe what the principle implies.
  5. Draw 2 graphs and label the x, y, and z axis on each. On 1 graph, draw the shape of the px orbital. On the other graph, draw the dx2-y2 orbital. Under each graph, write which angular momentum quantum number, l, each orbital corresponds to.
  6. Using complete subshell notation (1s22s22p6, and so forth), predict the electron configuration of each of the following atoms.
    1. N                     b. Sc                c. Na                d. As                e. Mn
  7. How many valence electrons do the following atoms have?
    1. P                      b. N                 c. K                 d. Ni                e. Ne
  8. In building the electron configuration of scandium, Sc, the last electron goes into which orbital, s, p, d, f, or g?
  9. Which of the following has 2 unpaired electrons?
    1. Mg       b. Si     c. S      d. Both Mg and S        e. Both Si and S
  10. What is the difference between an atom that is diamagnetic vs paramagnetic?
Answer #1

O [a) 767 nm = Red line (6) 399 nm = vided line * = C = 3163 (2) 7067X107 m wave length (2) 7= 3x108 = 3.9 x 1014 Hz (b) 3.99.2 En = -2.2410-18 _ _7667x10-19 NQ. n= 2.2010 V 762 n = 1.69 Å n = 0.629 xn 2 - 0-$29* (169)? M-51 i